Mechanics

A hollow smooth uniform sphere A of mass m rolls without sliding on a smooth horizontal surface. It collides elastically with another stationary smooth solid sphere B of same mass m and same radius. Find the ratio of kinetic energy of B to that of A just after the collision.

jay gohel
15 Years agoGrade

1 Answer

vikas askiitian expert
ApprovedApproved Tutor Answer15 Years ago

when two identical bodies collide & one body is at rest initially then 

linear momentam is totally transfered to other body...

let the velocity of sphere A initially was V....

final velocity of A= 0

final velocity of other sphere B is V ....                         (linear momentam is transfered)

kinetic energy of B =mV2/2 ..............1

kinetic energy of A = IW2/2

                           =mV2/3 ................2                (Isphere = 2/3 mr2 )

eq1 / eq2 =3/2